even A and B are such that P(A U B )=0.76,P(B)=0.4 Given that A and B are non mutually exclusive events that are independent calculate P(A)


1
Expert's answer
2021-06-16T06:21:28-0400

"A" and "B" are non mutually events: "P(A\cap B)\not=0."

"A" and "B" are independent events: "P(A\cap B)=P(A)P(B)."



"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=P(A)+P(B)-P(A)P(B)"

"P(A)=\dfrac{P(A\cup B)-P(B)}{1-P(B)}"

"=\dfrac{0.76-0.4}{1-0.4}=0.6"

"P(A)=0.6"